OpenStudy (anonymous):

LAST question promise If x-3, x+1, and 4x-2 are the first three terms of a geometric sequence, find x.

jimthompson5910 (jim_thompson5910):

Hint: If x-3, x+1, and 4x-2 are the first three terms of a geometric sequence then (4x-2)/(x+1) = (x+1)/(x-3) Solve that equation to find x

OpenStudy (anonymous):

THANKS

jimthompson5910 (jim_thompson5910):

yw, tell me what you get for x

OpenStudy (anonymous):

(4x-2)/(x+1) = (x+1)/(x-3) 4x-2=x-3 3x=-1 x=-1/3